# Compositional Design of Safety Controllers for Large-scale Stochastic Hybrid Systems

### v1

@@ Abstract.
In this work, we propose a compositional scheme based on small-gain reasoning for the safety controller synthesis of interconnected stochastic hybrid systems with both continuous evolutions and instantaneous jumps. In our proposed setting, we first offer an augmented scheme to represent each stochastic hybrid subsystem with continuous and discrete evolutions in a unified framework, ensuring that the state trajectories match those of the original hybrid systems. We then introduce the concept of augmented control sub-barrier certificates (A-CSBC) for each subsystem, which allows the construction of augmented control barrier certificates (A-CBC) for interconnected systems and their safety controllers under small-gain compositional conditions. We eventually leverage the constructed A-CBC and quantify a guaranteed probabilistic bound across the safety of the interconnected system. While the computational complexity of designing a barrier certificate and its safety controller grows polynomially with network dimension using sum-of-squares (SOS) optimization program, our compositional approach significantly reduces it to a linear scale with respect to the number of subsystems. We verify the efficacy of our proposed approach over an interconnected stochastic hybrid system composed of $1000$ nonlinear subsystems.
1. Introduction
@@
Stochastic hybrid systems (SHS) [BLE+06, CL06] are complex heterogeneous models, spanning a wide range of safety-critical applications, in which discrete dynamics describe computation and communication parts and continuous models capture physical components. Providing a formal verification and controller synthesis framework for this type of complex heterogeneous systems is naturally very challenging primarily due to (i) the stochastic nature of dynamics, (ii) the tight interaction between continuous and discrete components (in both space and time), (iii) the large dimension of state/input sets, and (iv) the need to manage complex logic requirements [BK08].
@@
Discretization-based approaches. In order to deal with the above-mentioned difficulties, one potential solution, proposed in the relevant literature, is to approximate original models by a simpler one with either a lower dimension (a.k.a., infinite abstractions) or with discrete-state sets (a.k.a., finite abstractions) [JP09, APLS08]. On the downside, the existing abstraction-based techniques rely on discretizing state/input sets and are not applicable to real-world high-dimensional systems due to the state-explosion problem. To make the abstraction-based techniques more scalable, compositional approaches have been proposed in the relevant literature that construct abstractions of large-scale systems based on those of smaller subsystems [LZ22, HHHK13, NZ20, LSAZ22, Lav24].
@@
Discretization-free approaches. To mitigate the computational complexity arising from abstraction-based approaches, control barrier certificates (CBC), a promising discretization-free approach, have garnered significant attention in recent years for the formal analysis of complex dynamical systems. Specifically, barrier certificates are Lyapunov-like functions whose initial level set separates an unsafe region from all system trajectories starting from a set of initial states. Consequently, the existence of such a function provides a (probabilistic) safety certificate over the system (cf. Fig. 1). This approach has been initially proposed by [PJ04, PJP07] and then widely employed for formal verification and synthesis of non-stochastic [BWAE15, ACE+19] and stochastic dynamical systems [ZRHH10, ASBA19, SDC19, Cla19, NZ22, LF24, NPNS24, ALZ22, JLZ22, WHL24], to name a few.
Figure 1. A control barrier certificate for stochastic dynamical systems. The initial level set of CBC is denoted by dashed line (i.e., $\mathcal{B}(x)=\mu$).
@@
Innovative findings. Despite the extensive above-mentioned literature on barrier certificate construction, there are no results available for large-scale stochastic hybrid systems. Our primary contribution is to propose a compositional scheme for the safety controller synthesis of interconnected SHS with both continuous evolutions, governed by stochastic differential equations with Brownian motions and Poisson processes, and instantaneous jumps, modeled by stochastic difference equations with additive noises. In our proposed setting, we first introduce an augmented framework to recast each stochastic hybrid subsystem with continuous and discrete evolutions via a unified framework, spanning both scenarios. We then introduce notions of augmented control sub-barrier certificates, computed for each subsystem, using which one can construct augmented control barrier certificates for interconnected systems together with their safety controllers under some small-gain compositional conditions. We eventually leverage the constructed augmented control barrier certificate and offer a certified probabilistic bound across the safety of the interconnected system. We verify the usefulness of our proposed approaches by applying them to an interconnected stochastic hybrid system consisting of $1000$ nonlinear subsystems. While the primary results of this work were initially presented in [LSF22], they are limited to monolithic systems and cannot handle models with more than six state dimensions due to the computational complexity involved in searching for control barrier certificates. In contrast, we propose here an innovative compositional technique to address large-scale SHS by constructing an augmented control barrier certificate for the interconnected system based on certificates of individual subsystems (cf. the $1000$-dimensional case study). While the computational complexity of designing a barrier certificate and its safety controller in [LSF22] grows polynomially with system dimension using sum-of-squares (SOS) optimization program, our compositional approach here significantly reduces it to a linear scale with respect to the number of subsystems. Additionally, we present the proofs for all statements that were not included in [LSF22].
Organization. The rest of the paper is structured as follows. Section 2 provides mathematical preliminaries and notations, and also the formal definition of (augmented) stochastic hybrid systems. In Section 3, we introduce notions of augmented control sub-barrier and barrier certificates for, respectively, augmented stochastic hybrid systems with and without disturbance signals. In Section 4, we propose a compositional approach based on small-gain reasoning to construct an augmented control barrier certificate for the interconnected system based on the certificates of individual subsystems. In Section 5, we provide required conditions for the construction of augmented control sub-barrier certificates based on control barrier certificates of original SHS. In Section 6, we employ sum-of-squares (SOS) optimization techniques to search for control barrier certificates of SHS. Section 7 verifies the proposed approaches over a nonlinear interconnected SHS, and concluding remarks are finally provided in Section 8.
2. Stochastic Hybrid Systems
2.1. Notation
@@
Sets of non-negative and positive integers are, respectively, denoted by :=$\mathbb{N}:=\{0,1,2,...\}$ and +=$\mathbb{N}^{+}=\{1,2,...\}$. Sets of real, positive and non-negative real numbers are represented by $\mathbb{R},\mathbb{R}^{+}$, and $\mathbb{R}_{0}^{+}$, respectively. Given $N$ vectors $x_{i}\in\mathbb{R}^{n_{i}}$, $x=[x_{1};...;x_{N}]$ represents the corresponding vector of dimension $\sum_{i}n_{i}$. Given a vector $x\in\mathbb{R}^{n}$ , we define the Hadamard power as $x^{\circ m}=[x_{1}^{m}~{}...~{}x_{n}^{m}]^{\top}$ for $m\in\mathbb{N}$. A (block) diagonal matrix in ${\mathbb{R}}^{N\times{N}}$ with diagonal matrix entries $(A_{1},\ldots,A_{N})$ and scalar entries $(a_{1},\ldots,a_{N})$ is denoted by $\mathsf{blkdiag}(A_{1},\ldots,A_{N})$ and $\mathsf{diag}(a_{1},\ldots,a_{N})$, respectively. We denote the Euclidean norm of a vector $x\in\mathbb{R}^{n}$ by $\|x\|$. Given sets $X_{i},i\in\{1,\ldots,N\}$, their Cartesian product is denoted by $\prod_{i=1}^{N}X_{i}$. Symbol $\mathds{I}_{n}$ denotes the identity matrix in $\mathbb{R}^{n\times{n}}$. Given a measurable function $f:\mathbb{N}\rightarrow\mathbb{R}^{n}$, the (essential) supremum of $f$ is denoted by $\|f\|_{\infty}:=\text{(ess)sup}\{\|f(k)\|,k\geq 0\}$.
2.2. Preliminaries
@@
2.3. Stochastic Hybrid Systems
In this work, our focus lies on stochastic hybrid systems, treating them as individual subsystems, as delineated in the following definition.
@@ 2.3. Stochastic Hybrid Systems
The SHS $\Phi_{i}$ can be characterized by the following stochastic differential and difference equations:
@@
$\Phi_{i}\!:\left\{\begin{array}[]{rl}\mathsf{d}x_{i}(t)\!=\!\!&f_{1_{i}}(x_{i}(t),\nu_{i}(t),w_{i}(t))\mathsf{d}t\!+\!\sigma_{i}(x_{i}(t))\mathsf{d}\mathbb{W}_{t}+\rho_{i}(x_{i}(t))\mathsf{d}\mathbb{P}_{t},~{}\quad\quad\quad\quad~{}~{}~{}\!\!\!\!t\in\mathbb{R}_{0}^{+}\backslash\Delta,\\ x_{i}(t)\!=\!\!&f_{2_{i}}(x_{i}(t^{-}),\nu_{i}(t),w_{i}(t),\varsigma_{i}(t)),\hskip 144.54004pt\!\!\!t\in\Delta,\end{array}\right.$ (2.3)
@@
where $\Delta=\{t_{k}\}_{k\in\mathbb{N}}$ with $t_{k+1}-t_{k}\in\{\varepsilon_{1_{i}}\tau,\ldots,\varepsilon_{2_{i}}\tau\}$ for jump parameters $\tau\in\mathbb{R}^{+}$ and $\varepsilon_{1_{i}},\varepsilon_{2_{i}}\in\mathbb{N}^{+}$, $\varepsilon_{1_{i}}\leq\varepsilon_{2_{i}}$. Moreover, $f_{2_{i}}(x_{i}(t^{-}))=\lim_{t\to t_{k}^{-}}f_{2_{i}}(x_{i}(t))$, indicating the value of $f_{2_{i}}(x_{i}(t))$ as $t$ goes to $t_{k}$ from left. Furthermore, $x_{i}:\mathbb{R}_{0}^{+}\rightarrow\mathbb{R}^{n_{i}}$ is the state signal (right-continuous for all $t\in\mathbb{R}_{0}^{+}$) and $\nu_{i}(\cdot)\in\mathcal{U}_{i},w_{i}(\cdot)\in\mathcal{W}_{i}$ are input and disturbance signals. The random sequence $x_{x_{0_{i}}\nu_{i}w_{i}}(t)$ fulfilling (2.3) is called the state trajectory (a.k.a. solution) of $\Phi_{i}$ at time $t\in\mathbb{R}_{0}^{+}$ under signals $\nu_{i}(\cdot),w_{i}(\cdot)$ starting from $x_{0_{i}}$.
@@
Given that control inputs in practice are often fed by a digital controller, we are interested here in sampled-data hybrid systems as
$\mathcal{U}_{\tau_{i}}\!\!=\!\Big{\{}\nu_{i}\!:\mathbb{R}_{0}^{+}\!\rightarrow\!U_{i}\,\big{|}\,\nu_{i}(t)\!=\!\nu_{i}(t_{k-1}),t\!\in\![t_{k-1},t_{k}],k\!\in\!\mathbb{N}\Big{\}},$ (2.4)
@@
2.4. Interconnected SHS

### v2

@@ Abstract
In this work, we propose a compositional scheme based on small-gain reasoning for the safety controller synthesis of interconnected stochastic hybrid systems with both continuous evolutions and instantaneous jumps. In our proposed setting, we first offer an augmented scheme to represent each stochastic hybrid subsystem with continuous and discrete evolutions in a unified framework, ensuring that the state trajectories match those of the original hybrid systems. We then introduce the concept of augmented control sub-barrier certificates (A-CSBC) for each subsystem, which allows the construction of augmented control barrier certificates (A-CBC) for interconnected systems and their safety controllers under small-gain compositional conditions. We eventually leverage the constructed A-CBC and quantify a guaranteed probabilistic bound across the safety of the interconnected system. While the computational complexity of designing a barrier certificate and its safety controller grows polynomially with network dimension using sum-of-squares (SOS) optimization program, our compositional approach significantly reduces it to a linear scale with respect to the number of subsystems. We verify the efficacy of our proposed approach over an interconnected stochastic hybrid system composed of $1000$ nonlinear subsystems with two distinct interconnection topologies.
keywords:
Stochastic hybrid systems, scalable safety controllers, control barrier certificates, small-gain reasoning
, , , and
1 Introduction
@@
Stochastic hybrid systems (SHS) (Blom et al., 2006; Cassandras & Lygeros, 2006) are complex heterogeneous models, spanning a wide range of safety-critical applications, in which discrete dynamics describe computation and communication parts and continuous models capture physical components. Providing a formal verification and controller synthesis framework for such complex heterogeneous systems is naturally very challenging primarily due to (i) the stochastic nature of dynamics, (ii) the tight interaction between continuous and discrete components (in both space and time), (iii) the large dimension of state/input sets, and (iv) the need to manage complex logic requirements (Baier & Katoen, 2008).
@@
Discretization-based approaches. In order to deal with the above-mentioned difficulties, one potential solution, proposed in the relevant literature, is to approximate original models by a simpler one with either a lower dimension (a.k.a., infinite abstractions) or with discrete-state sets (a.k.a., finite abstractions) (Julius & Pappas, 2009; Abate et al., 2008). On the downside, the existing abstraction-based techniques rely on discretizing state/input sets and are not applicable to real-world high-dimensional systems due to the state-explosion problem. To make the abstraction-based techniques more scalable, compositional approaches have been proposed in the relevant literature that construct abstractions of large-scale systems based on those of smaller subsystems (Lavaei & Zamani, 2022; Hahn et al., 2013; Nejati et al., 2020; Lavaei, Soudjani, Abate & Zamani, 2022; Lavaei, 2019).
@@
Discretization-free approaches. To mitigate the computational complexity arising from abstraction-based approaches, control barrier certificates (CBC), a promising discretization-free approach, have garnered significant attention in recent years for the formal analysis of complex dynamical systems. Specifically, barrier certificates are Lyapunov-like functions whose initial level set separates an unsafe region from all system trajectories starting from a set of initial states. Consequently, the existence of such a function provides a (probabilistic) safety certificate over the system. This approach has been initially proposed by Prajna & Jadbabaie (2004); Prajna et al. (2007) and then widely employed for formal verification and synthesis of non-stochastic (Borrmann et al., 2015; Ames et al., 2019; Wooding et al., 2025) and stochastic dynamical systems (Zhang et al., 2010; Jahanshahi et al., 2022; Ahmadi et al., 2019; Nejati et al., 2024; Santoyo et al., 2019; Clark, 2019; Nejati & Zamani, 2022; Lavaei & Frazzoli, 2024; Nejati & Schmuck, 2024), to name a few.
@@
Innovative findings. Despite the extensive above-mentioned literature on barrier certificate construction, there are no results available for large-scale stochastic hybrid systems. Our primary contribution is to propose a compositional scheme for the safety controller synthesis of interconnected SHS with both continuous evolutions, governed by stochastic differential equations with Brownian motions and Poisson processes, and instantaneous jumps, modeled by discrete-time stochastic reset equations with additive noises. In our proposed setting, we first introduce an augmented framework to recast each stochastic hybrid subsystem with continuous and discrete evolutions via a unified framework, spanning both scenarios. We then introduce notions of augmented control sub-barrier certificates, computed for each subsystem, using which one can construct augmented control barrier certificates for interconnected systems together with their safety controllers under some small-gain compositional conditions. We eventually leverage the constructed augmented control barrier certificate and offer a certified probabilistic bound across the safety of the interconnected system.
While a recent work by Lavaei & Frazzoli (2024) addresses networks of discrete-time switched systems with jumps acting solely through the transition map, our work considers a broader class of stochastic hybrid systems. These include both continuous- and discrete-time (reset) dynamics, with switching in both flow and jump components, modeled via stochastic differential and reset equations driven by Brownian motion, Poisson processes, and additive noise. This richer structure requires safety analysis for both evolution modes and necessitates additional design elements (see Theorem 5.4) to ensure soundness across the hybrid dynamics. Moreover, unlike (Lavaei & Frazzoli, 2024), which lacks control inputs and focuses on switching signal design under dwell-time constraints, our model explicitly incorporates control inputs in both the flow and jump dynamics. Accordingly, the controller synthesis problem becomes more challenging in our setting, as it involves two distinct types of adversarial inputs: (i) disturbance inputs, which capture the influence of other subsystems, and (ii) switching between flow and jump modes in a hybrid setting with both continuous- and discrete-time dynamics. Finally, while Lavaei & Frazzoli (2024) adopts a dissipativity-based compositional method, our approach is grounded in small-gain reasoning, which demands carefully designing the disturbance gains to be sufficiently small to enable lifting safety guarantees from individual subsystems to the entire interconnected network. This is particularly challenging in stochastic hybrid systems due to their hybrid nature.
While the primary results of this work were initially presented in (Lavaei, Soudjani & Frazzoli, 2022), they are limited to monolithic systems and cannot handle models with more than six state dimensions due to the computational complexity involved in searching for control barrier certificates. In contrast, we propose here an innovative compositional technique to address large-scale SHS by constructing an augmented control barrier certificate for the interconnected system based on certificates of individual subsystems (cf. the $1000$-dimensional case study). While the computational complexity of designing a barrier certificate and its safety controller in (Lavaei, Soudjani & Frazzoli, 2022) grows polynomially with system dimension using sum-of-squares (SOS) optimization program, our compositional approach here significantly reduces it to a linear scale with respect to the number of subsystems (cf. Subsection 7.4). Moreover, a key challenge in analyzing networked stochastic hybrid systems lies not only in the high dimensionality inherent to such systems, but also in the coupling between subsystems. This coupling, modeled as the adversarial input in each subsystem’s dynamics (see (2.4)), poses a central difficulty specifically due to the hybrid nature of the system, necessitating a more careful treatment of these interconnections (cf. Remark 5.6), which is absent in (Lavaei, Soudjani & Frazzoli, 2022). Additionally, we present the proofs for all statements that were not included in (Lavaei, Soudjani & Frazzoli, 2022).
2 Stochastic Hybrid Systems
2.1 Notation
@@
Sets of non-negative and positive integers are, respectively, denoted by $\mathbb{N}:=\{0,1,2,$...\} and $\mathbb{N}^{+}=\{1,2,$...\}. Sets of real, positive and non-negative real numbers are represented by $\mathbb{R},\mathbb{R}^{+}$, and $\mathbb{R}_{0}^{+}$, respectively. Given $N$ vectors $_{i}\in\mathbb{R}^{n_{i}}, x=[x$$x_{1};...;x_{N}$] represents the corresponding vector of dimension $\sum_{i}n_{i}$. Given a vector $x\in\mathbb{R}^{n}$ , we define the Hadamard power as $x^{\circ m}=[x_{1}^{m}~...~x_{n}^{m}]^{\top}$ for $m\in\mathbb{N}$. A (block) diagonal matrix in ${\mathbb{R}}^{N\times{N}}$ with diagonal matrix entries $(A_{1},\ldots,A_{N})$ and scalar entries $(a_{1},\ldots,a_{N})$ is denoted by $\mathsf{blkdiag}(A_{1},\ldots,A_{N})$ and $\mathsf{diag}(a_{1},\ldots,a_{N})$, respectively. The notation $\{a_{ij}\}$ denotes a matrix whose entry in the $i$-th row and $j$-th column is given by $a_{ij}$. We denote the Euclidean norm of a vector $x\in\mathbb{R}^{n}$ by $\|x\|$. Given sets $X_{i},i\in\{1,\ldots,N\}$, their Cartesian product is denoted by $\prod_{i=1}^{N}X_{i}$. The empty set is represented by $\emptyset$. Symbol $\mathds{I}_{n}$ denotes the identity matrix in $\mathbb{R}^{n\times{n}}$. Given a measurable function $f:\mathbb{N}\rightarrow\mathbb{R}^{n}$, the (essential) supremum of $f$ is denoted by $\|f\|_{\infty}:=\text{(ess)sup}\{\|f(k)\|,k\geq 0\}$.
2.2 Preliminaries
@@
2.3 Stochastic Hybrid Systems
In this work, our focus lies on stochastic hybrid systems, treating them as individual subsystems. We begin by formally introducing the notion of a stochastic hybrid system (SHS), which generalizes the impulsive system concept in (Swikir et al., 2020, Definition 1) by including different sources of noises and disturbances.
@@ 2.3 Stochastic Hybrid Systems
The SHS $\Phi_{i}$ can be characterized by the following stochastic differential and reset equations:
@@
$\Phi_{i}\!:\left\{\hskip-4.2679pt\begin{array}[]{rl}\mathsf{d}x_{i}(t)\!=\!\!&f_{1_{i}}(x_{i}(t),\nu_{i}(t),w_{i}(t))\mathsf{d}t\!+\!\sigma_{i}(x_{i}(t))\mathsf{d}\mathbb{W}_{t}\\ &+\rho_{i}(x_{i}(t))\mathsf{d}\mathbb{P}_{t},~\quad\quad\quad\quad\quad\quad~~~t\in\mathbb{R}_{0}^{+}\backslash\Delta,\\ x_{i}(t)\!=\!\!&{f_{2_{i}}(x_{i}(t^{-}),\nu_{i}(t),w_{i}(t^{-}),\varsigma_{i}(t))},\quad\!\!\!t\in\Delta,\end{array}\right$. (2.4)
@@
where $\Delta=\{t_{k}\}_{k\in\mathbb{N}}$ with $t_{k+1}-t_{k}\in\{\varepsilon_{1_{i}}\tau,\ldots,\varepsilon_{2_{i}}\tau\}$ for jump parameters $\tau\in\mathbb{R}^{+}$ and $\varepsilon_{1_{i}},\varepsilon_{2_{i}}\in\mathbb{N}^{+}$, $\varepsilon_{1_{i}}\leq\varepsilon_{2_{i}}$. Here, $\mathbb{W}_{t}$ denotes the Brownian motion, and $\mathbb{P}_{t}$ represents the Poisson process with rate $\lambda$. Moreover, $f_{2_{i}}(x_{i}(t^{-}),\nu_{i}(t),w_{i}(t^{-}),\varsigma_{i}(t))=\lim_{t\to t_{k}^{-}}f_{2_{i}}(x_{i}(t),\nu_{i}(t),w_{i}(t),\varsigma_{i}(t))$, indicating the value of $f_{2_{i}}(x_{i}(t),\nu_{i}(t),w_{i}(t),\varsigma_{i}(t))$ as $t$ goes to $t_{k}$ from left. Furthermore, $x_{i}:\mathbb{R}_{0}^{+}\rightarrow\mathbb{R}^{n_{i}}$ is the state signal (right-continuous for all $t\in\mathbb{R}_{0}^{+}$) and $\nu_{i}(\cdot)\in\mathcal{U}_{i},w_{i}(\cdot)\in\mathcal{W}_{i}$ are input and disturbance signals. The random sequence $x_{x_{0_{i}}\nu_{i}w_{i}}(t)$ fulfilling (2.4) is called the state trajectory (a.k.a. solution) of $\Phi_{i}$ at time $t\in\mathbb{R}_{0}^{+}$ under signals $\nu_{i}(\cdot),w_{i}(\cdot)$ starting from $x_{0_{i}}$.
Remark 2.2.
In the SHS defined in Definition 2.1, jumps at time instants $t\in\Delta$ introduce discontinuities in the state $x_{i}(t)$, which evolves discretely at such times. To ensure the well-posedness of the model, it is necessary to specify the left-hand limit $x_{i}(t^{-})$ at each jump instant. The same requirement applies to the adversarial inputs $w_{i}(t)$, which represent the (possibly discontinuous) states of neighboring subsystems (cf. (2.7)). In contrast, the control input $\nu_{i}(t)$ is user-designed and can be applied as a continuous-time signal without needing to exhibit jumps (Swikir et al., 2020). The additive noise $\varsigma_{i}(t)$ is also treated as externally imposed and does not require specification at $t^{-}$ as it is not necessarily exhibiting jumps.
Remark 2.3.
The drift, diffusion, reset functions, and the transition map of the stochastic hybrid system are all assumed to be globally Lipschitz continuous. This assumption, stated in Definition 2.1, ensures the existence, uniqueness, and strong Markov property of the solution process $x_{x_{0_{i}}\nu_{i}w_{i}}(t)$, and thereby, the well-posedness of the stochastic hybrid system. Of note is, since our CBC analysis is restricted to compact domains, any continuous function defined on such domains is automatically globally Lipschitz continuous.
To further clarify the modeling capabilities of such systems and the roles of the Brownian motion and Poisson process, we present the following illustrative example.
Example 2.4.
Consider a network of $N$ rooms, each equipped with a local heater. The dynamics in (2.4) can model each room, where $x_{i}(t)$ denotes its temperature, controlled by the heater control input $\nu_{i}(t)$, and influenced by $w_{i}(t)$, the inter-room heat transfer (e.g., through shared walls). The Brownian motion $\mathbb{W}_{t}$ models continuous environmental uncertainties (e.g., brief door/window openings), while the Poisson process $\mathbb{P}_{t}$ captures discrete events like people entering, causing sudden temperature changes. The jump times $\Delta=\{t_{k}\}_{k\in\mathbb{N}}$ correspond to sudden thermostat reconfigurations (e.g., switching between ECO and COMFORT modes), occurring at intervals constrained by $\varepsilon_{1_{i}}\tau\leq t_{k+1}-t_{k}\leq\varepsilon_{2_{i}}\tau$.
@@
Given that control inputs in practice are often fed by a digital controller, we are interested here in sampled-data hybrid systems in which the control inputs are considered to be piecewise constant as
$\mathcal{U}_{\tau_{i}}\!\!=\!\Big\{\nu_{i}\!:\mathbb{R}_{0}^{+}\!\rightarrow\!U_{i}\,\big|\,\nu_{i}(t)\!=\!\nu_{i}(t_{k-1}),{t\!\in\![t_{k-1},t_{k}),k\!\in\!\mathbb{N}^{+}}\!\!\Big\},$ (2.5)
@@
Remark 2.5.
While our theoretical framework does not require the control inputs $\nu_{i}$ to be piecewise constant, we adopt this assumption as control inputs in digital practical implementations are typically applied using a zero-order hold mechanism, which naturally results in piecewise constant signals.
2.4 Interconnected SHS